OpenStudy (anonymous):

Write the equation in slope-intercept form. 3x + y = 16

OpenStudy (anonymous):

ok easy stuff Slope Intercept Form= y=mx+b where mx= the slope and integer is the y intercept Can you identify the slope and Y intercept

OpenStudy (kaiz122):

the slope intercept form is y=mx+b, where m is the slope

OpenStudy (anonymous):

No o:

OpenStudy (anonymous):

hint: the slope is the the number that includes x

OpenStudy (anonymous):

you there

OpenStudy (anonymous):

i dont get this

OpenStudy (anonymous):

okay look mx= 3x and y intercept=16 and you have to put your equation is y=mx+b form